# HybridoNet-Adapt: A Domain-Adapted Framework for Accurate Lithium-Ion Battery RUL Prediction

### v1

@@ Abstract
Accurate prediction of the remaining useful life (RUL) in Lithium-ion battery (LIB) health management systems is crucial for ensuring reliability and safety. Current methods typically assume that training and testing data share the same distribution, overlooking the benefits of incorporating diverse data sources to enhance model performance. To address this limitation, we introduce a data-independent RUL prediction framework along with its domain adaptation (DA) approach, which leverages heterogeneous data sources for improved target predictions. Our approach integrates comprehensive data preprocessing—comprising feature extraction, denoising, and normalization—with a data-independent prediction model that combines Long Short-Term Memory (LSTM), Multihead Attention, and a Neural Ordinary Differential Equation (NODE) block, termed HybridoNet. The domain-adapted version, HybridoNet-Adapt, is trained using a novel technique inspired by the Domain-Adversarial Neural Network (DANN) framework, a regression ensemble method, and Maximum Mean Discrepancy (MMD) to learn domain-invariant features from labeled cycling data in the source and target domain. Experimental results demonstrate that our approach outperforms state-of-the-art techniques, providing reliable RUL predictions for real-world applications.
Keywords—Battery Health Management, Lithium-ion Batteries, Remaining Useful Life, Signal Processing, Data-Independent Prediction
journal: Mechanical Systems and Signal Processing
@@
[1]organization=AIWARE Limited Company, addressline=17 Huynh Man Dat Street, Hoa Cuong Bac Ward, Hai Chau District, city=Danang, citysep=, postcode=550000, country=Vietnam
@@
[2]organization=AIT Austrian Institute of Technology GmbH, addressline=Giefinggasse 4, city=Vienna, citysep=, postcode=1210, country=Austria
@@
[3]organization=Software Engineering Department - FPT University, city=Da Nang, postcode=550000, country=Vietnam
@@ 1.1 Motivations
Lithium-ion batteries (LIBs) [11], renowned for their affordability and high energy density, are extensively utilized [14, 8, 50, 23] in electric vehicles (EVs), mobile phones, and energy storage stations. The global lithium-ion battery (LIB) market is projected to surpass 170 billion dollars by 2030 [33]. With the wide-ranging adoption of LIBs, interest in battery health management (BHM) has surged within both academia and industry in recent years. In a BHM system, several common and essential techniques are employed, including thermal management [21, 49], fault diagnosis/detection [7], state of charge (SOC) and state of health (SOH) estimation [28], RUL prediction [24, 31], and cycle life early prediction [38, 16, 29]. In the lifespan of a LIB, RUL prediction aims to forecast the remaining life, which can be measured by capacity or the number of remaining cycles the battery cell can undergo before reaching the end of life (EOL). RUL prediction methods fall into three categories: model-based, data-driven, and hybrid approaches.
Traditional model-based approaches often utilize physics-based degradation models, such as the Double Exponential Model (DEM) [29], two-phase degradation models [42], and Markov Models [51], constructed using early-cycle data (200-500 cycles). These models aim to forecast the entire battery’s capacity degradation curve. However, relying solely on maximum discharge capacity degradation and early-cycle data often leads to inaccuracies due to the influence of various factors (current, voltage, temperature, time) and sudden changes in degradation trends [29, 42].
Data-driven models predict the RUL of LIBs by analyzing present cycling data. Techniques like dual-input Deep Neural Networks (DNN) [46], 1D Convolutional Neural Networks (1DCNN) [22], Dense layers [19], Long Short-Term Memory (LSTM) networks [13], and Echo State Networks (ESN) [18] have shown superior performance. These methods require large amounts of high-quality data.
@@
Accurately predicting the RUL of LIBs remains challenging due to their nonlinear degradation processes and the need for highly accurate prediction models. The specific challenges will be discussed in the next section.
@@ 1.2 Problem Statement
As summarized in Table 1, current state-of-the-art studies categorize RUL prediction methods into two primary approaches: historical data-independent methods, which estimate the current RUL based on present cycling data and a few preceding cycles, and historical data-dependent methods, which leverage extensive early-cycle data to predict the battery’s full lifespan.
@@
Many existing methods focus on predicting RUL based on estimated maximum discharge capacity thresholds. However, accurately forecasting the remaining cycles provides clearer insights into the battery’s usable life.
Datasets such as the TRI dataset (124 cells, fast-charging) and the LHP dataset (77 cells, diverse discharge) provide extensive charge-discharge scenarios, making them well-suited for both training and validation of data-driven models. In contrast, smaller datasets like the Oxford Battery dataset (13 cells) and NASA battery datasets (4–34 cells) limit model robustness in real-world failure prediction.
Signal preprocessing techniques are generally categorized into statistical feature extraction methods, such as mean and standard deviation (Std), and variational decomposition methods like Empirical Mode Decomposition (EMD) and Variational Mode Decomposition (VMD). While variational decomposition increases the complexity of the input signal and often maintains or enlarges its original size, statistical feature extraction significantly reduces the signal’s dimensionality and effectively captures key signal characteristics, making it well-suited for real-time applications and high-accuracy predictions in industrial settings.
Model-based and hybrid approaches typically rely on early-cycle data for RUL prediction, yet each battery exhibits unique degradation patterns over its lifespan requiring adaptive data-driven strategies. Moreover, Transfer learning methods such as domain-adversarial neural networks (DANN) [10, 47] and Generative Adversarial Networks (GANs) [12] offer effective solutions for adapting to other source’s degradation patterns to improve RUL prediction in the target domain. To address these challenges, our proposed approach will be represented in the next section.
@@ 1.3 Main Contribution
We propose a historical data-independent method that leverages only present and recent cycling data, eliminating the need for early-cycle data while ensuring accurate RUL forecasts. Additionally, we conduct large-scale experiments on two largest available datasets of A123 APR18650M1A cells [38, 46], covering diverse charging and discharging conditions. These evaluations validate the effectiveness and real-world applicability of our data-driven approach. Our key contributions include:
  1. 1.
    A RUL Prediction Module: We propose a comprehensive RUL prediction framework that consists of a signal processing phase and a regression neural network, named HybridoNet. The signal processing phase extracts key degradation features from raw battery signals, while HybridoNet, composed of LSTM, Multihead Attention [41], and Neural Ordinary Differential Equation (NODE) [4] blocks. This data-driven approach achieves superior accuracy compared to existing methods.
  2. 2.
    Enhancing Neural Networks with Domain Adaptation: To further improve generalization across different battery degradation conditions, we introduce HybridoNet-Adapt, an extension of HybridoNet incorporating a Domain Adaptation (DA) technique inspired by DANN [10]. This adaptation mechanism helps the model learn transferable representations, mitigating discrepancies between different operating conditions and improving RUL prediction in real-world applications.
@@ 2 Preliminaries
This section provides an overview of HybridoNet’s key components, including LSTM, Multihead Attention, and NODE. It also represents the DA-based strategy that inspires HybridoNet-Adapt.
2.1 LSTM:
A recurrent neural network (RNN) [35] architecture designed to overcome the vanishing gradient problem by introducing gating mechanisms. Its operations are defined by:
@@ 2.2 Multihead Attention:
2.3 NODE:
A framework that extends deep learning architectures to model continuous dynamics. In NODE, the evolution of a hidden state $h(t)$ is governed by an ordinary differential equation:
@@ 2.3 NODE:
where $f$ is a neural network parameterized by $\theta$. The final state is obtained by solving this ODE over a time interval, which provides a flexible and memory-efficient representation.
2.4 DANN:
A framework facilitates domain adaptation by learning domain-invariant features. It encourages the feature extractor $G_{F}$ to generate features that are indistinguishable across domains. The overall objective is formulated as:
$\min_{G_{F},G_{Y}}\max_{G_{D}}\ \mathcal{L}(G_{F},G_{Y},G_{D})$ $=\mathbb{E}_{(x,y)\sim S}\left[\mathcal{L}_{y}(G_{Y}(G_{F}(x)),y)\right]$
$\quad-\lambda\mathbb{E}_{x\sim S\cup T}\left[\mathcal{L}_{d}(G_{D}(G_{F}(x)),d)\right],$
where $G_{Y}$ is the label predictor, $G_{D}$ is the domain discriminator, $S$ and $T$ denote the source and target domains, $d$ represents domain labels, and $\lambda$ balances classification and domain confusion losses.
@@ 3.1 Overall architecture
Figure 1: Overall architecture of our RUL prediction process for Lithium-ion battery cells.
@@
The signal preprocessing phase begins with applying a median filter [20] to smooth out sudden peaks in the raw signal vectors. Each signal vector—voltage, current, and capacity—is then processed through feature extraction methods, including mean, standard deviation (Std), minimum (Min), maximum (Max), variance (Var), and median (Med) [9, 39]. The representation of data for each cycle is given by $X^{i}=\left[x_{\text{current}}^{i},x_{\text{voltage}}^{i},x_{\text{capacity}}^{i}\right],\quad X^{i}\in\mathbb{R}^{3\times 6},$ where 3 corresponds to the three signal types (voltage, current, and capacity), and 6 represents the extracted features (Mean, Std, Min, Max, Var, Med). Each input sample for the RUL prediction model consists of 10 selected cycles sampled from a 30-cycle window (one cycle every three cycles) [30]. The input sample can be represented as $\mathbf{X^{i}}=\left[X^{1},X^{2},\dots,X^{10}\right],\quad\mathbf{X}\in\mathbb{R}^{10\times 3\times 6}.$ The output’s shape of input data after feature extraction step is $\mathbf{X}\in\mathbb{R}^{N\times 10\times 3\times 6},$ where $N$ denotes the number of samples. In the normalization step, MinMaxScaler is applied. $\mathbf{X}$ is first reshaped from $\mathbb{R}^{N\times 10\times 3\times 6}$ to $\mathbb{R}^{(N\times 10)\times 18}$, where 18 represents the flattened $6\times 3$ feature matrix. MinMaxScaler is then fitted and applied to scale the data between 0 and 1. Finally, the normalized data is reshaped back to $\mathbb{R}^{N\times 10\times 3\times 6}$ for further processing.
In the prediction phase, the RUL prediction model, HybridoNet-Adapt, consisting of multiple blocks, is used to map $\mathbf{X}$ to the predicted RUL $\mathbf{Y}\in\mathbb{R}^{N\times 1}$. The details of the proposed RUL prediction models are presented in the following sections.
3.2 HybridoNet: A Proposed RUL model
HybridoNet consists of two key components: a predictor $G^{T}_{Y}$ and a feature extractor $G_{F}$ module. The feature extractor integrates an LSTM, a Multihead Attention mechanism, and a NODE block. In the NODE block, a linear layer is employed to approximate the functionality of multiple linear layers, thereby reducing computational overhead. This block applies a continuous transformation function: $\frac{dh(t)}{dt}=f(h(t),t,\theta),$ where $t$ represents the time steps, $\theta$ denotes the trainable parameters of $f$, and $f$ is a trainable function. In our NODE block, $f$ is implemented as a linear layer. The initial condition of the NODE block is given by $h(t_{0})$. Solving the ODE over the interval $[t_{0},t_{1}]$ produces the transformed state $h(t_{1})$, which serves as the NODE block output:
@@ 3.2 HybridoNet: A Proposed RUL model
Where, $t_{0}$ and $t_{1}$ are set to 0 and 1, respectively, based on experimental findings (see Figure 7).
The predictor module consists of three linear layers, each followed by ReLU [5] activation, 1D Batch Normalization [17], and Dropout [15], respectively, with a sigmoid activation [48] layer at the end. The HybridoNet model is trained using the Mean Squared Error (MSE) [44] loss function to optimize the model via an optimizer. The MSE loss is defined as:
$\mathcal{L}_{\text{MSE}}=\frac{1}{N}\sum_{i=1}^{N}(G^{S}_{Y}(G_{F}(X_{i}))-Y_{i})^{2}$
where $Y_{i}$ is the RUL ground truth. Note that the proposed HybridoNet is not the main focus of this paper. Instead, we primarily introduce HybridoNet-Adapt, a DA-based extension of HybridoNet, which will be presented in the next section.
3.3 HybridoNet-Adapt: An Enhanced Version Leveraging Domain Adaptation
Figure 2: The training process for the proposed HybridoNet-Adapt.
As shown in Figure 2, HybridoNet-Adapt extends HybridoNet with a DA technique. It includes a source predictor $G^{S}_{Y}$, a target predictor $G^{T}_{Y}$, and a feature extractor $G_{F}$, all sharing the same architecture as their counterparts in HybridoNet (Section 3.2). The target and source prediction in HybridoNet-Adapt are computed as:
@@ 3.3 HybridoNet-Adapt: An Enhanced Version Leveraging Domain Adaptation
$\hat{Y}^{S}_{i}(X_{i})$ $=G^{S}_{Y}\Big{(}G_{F}(X_{i})\Big{)}.$ (2)
Here, $\theta^{S}$ and $\theta^{T}$ are trainable trade-off parameters that balance the contributions of the source and target predictors, and $\hat{Y}_{i}$ represents the predicted output.
The HybridoNet-Adapt is optimized by two losses: the mean squared error (MSE) loss, $\mathcal{L}_{\text{MSE}}$, for the regression outputs, and the maximum mean discrepancy (MMD) loss, $\mathcal{L}_{\text{MMD}}$, to align feature embeddings. The total loss is defined as follows:
@@
$+\lambda\,\mathcal{L}_{\text{MMD}}\Big{(}G_{F}(X^{S}_{i}),G_{F}(X^{T}_{i})\Big{)}.$
Here, $X^{S}_{i}$ and $X^{T}_{i}$ represent input samples from the source and target domains, respectively, and $Y^{S}_{i}$ and $Y^{T}_{i}$ denote the corresponding RUL ground truth values.
The MMD loss measures the discrepancy between the distributions of feature embeddings. Given feature sets $\{X^{S}_{i}\}_{i=1}^{n}$ from the source domain and $\{X^{T}_{i}\}_{i=1}^{m}$ from the target domain, it is defined as:
$\mathcal{L}_{\text{MMD}}=\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}k(X^{S}_{i},X^{S}_{j})+\frac{1}{m^{2}}\sum_{i=1}^{m}\sum_{j=1}^{m}k(X^{T}_{i},X^{T}_{j})-\frac{2}{nm}\sum_{i=1}^{n}\sum_{j=1}^{m}k(X^{S}_{i},X^{T}_{j}),$
where $k(\cdot,\cdot)$ is a kernel function, commonly chosen as the Gaussian kernel: $k(x,y)=\exp\left(-\frac{\|x-y\|^{2}}{2\sigma^{2}}\right),$ with $\sigma$ as the kernel bandwidth parameter.
The hidden dimension of the LSTM and Multihead Attention is set to 128, with 2 LSTM layers. The hidden dimensions of the linear layers in the predictor are $[128,64,32,1]$, with a dropout rate of 0.1. These values are determined based on the experimental results presented in the next section.
In the next section, multiple experiments are conducted to determine the optimal configuration of HybridoNet-Adapt, and demonstrating the superiority over state-of-the-art methods.

### v2

@@ Abstract
Accurate prediction of the Remaining Useful Life (RUL) in Lithium-ion battery (LIB) health management systems is essential for ensuring operational reliability and safety. However, many existing methods assume that training and testing data follow the same distribution, limiting their ability to learn domain-invariant features and generalize effectively to unseen target domains. To approach this opportunity, we propose a novel RUL prediction framework and incorporates an domain adaptation (DA) technique. Our framework integrates a comprehensive signal preprocessing pipeline—including noise-reduction, feature extraction, and normalization—with a robust deep learning model named HybridoNet-Adapt. The model comprises a feature extraction block built from Long Short-Term Memory (LSTM), Multihead Attention, and Neural Ordinary Differential Equation (NODE) layers, followed by two predictor modules implemented using linear layers. These predictors are balanced using trainable trade-off parameters. To enhance generalization to unseen data distributions, we introduce a domain adaptation strategy inspired by the Domain-Adversarial Neural Network (DANN) framework, replacing adversarial loss with Maximum Mean Discrepancy (MMD) to learn domain-invariant features. This approach enables effective transfer learning from source to target domains using cycling data. Experimental results demonstrate that our proposed method significantly outperforms traditional machine learning models such as XGBoost and Elastic Net, as well as deep learning baselines like Dual-input DNN, in predicting RUL. These findings highlight the potential of HybridoNet-Adapt for reliable and scalable Battery Health Management (BHM).
Keywords—Battery Health Management, Lithium-ion Batteries, Remaining Useful Life, Signal Processing, Domain Adaptation
@@
[1]organization=AIWARE Limited Company, addressline=17 Huynh Man Dat Street, Hoa Cuong Bac Ward, Hai Chau District, city=Da Nang, postcode=550000, country=Vietnam
@@
[2]organization=AIT Austrian Institute of Technology GmbH, addressline=Giefinggasse 4, city=Vienna, postcode=1210, country=Austria
@@
[3]organization=Software Engineering Department, FPT University, city=Da Nang, postcode=550000, country=Vietnam
[…]
[4]organization=Faculty of Information Technology, Ton Duc Thang University, city=Ho Chi Minh City, postcode=700000, country=Vietnam
[…]
[5]organization=Laboratory of Environmental Sciences and Climate Change, Institute for Computational Science and Artificial Intelligence, Van Lang University, city=Ho Chi Minh City, postcode=700000, country=Vietnam
[…]
[6]organization=Faculty of Environment, School of Technology, Van Lang University, city=Ho Chi Minh City, postcode=700000, country=Vietnam
@@ 1.1 Motivations
Lithium-ion batteries (LIBs) [15], renowned for their affordability and high energy density, are extensively utilized [18, 11, 55, 28] in electric vehicles (EVs), portable devices, and energy storage stations. The global lithium-ion battery (LIB) market is projected to surpass 170 billion dollars by 2030 [40]. With the wide-ranging adoption of LIBs, interest in battery health management (BHM) has surged within both academia and industry in recent years. In a BHM system, several common and essential techniques are employed, including thermal management [26, 54], fault diagnosis/detection [10], state of charge (SOC) [37, 2, 20] and state of health (SOH) [34] estimation, remaining useful life (RUL) prediction [29, 38], and cycle life early prediction [44, 21, 31, 57]. Among these, RUL prediction plays a crucial role in ensuring the reliability, safety, and optimal performance of LIBs over their lifespan. RUL can be assessed based on a capacity metric, such as SOH, or the number of remaining cycles the battery can undergo before reaching its end of life. Accurate RUL prediction allows for proactive maintenance, minimizing downtime, and enhancing the operational efficiency of LIBs. RUL prediction falls into three categories: model-based, data-driven, and hybrid approaches.
Traditional model-based approaches often utilize physics-based degradation models, such as the Double Exponential Model (DEM) [35], two-phase degradation models [48], and Markov Models [56], constructed using early-cycle data (200-500 cycles). These models aim to forecast the entire battery’s capacity degradation curve. However, relying solely on maximum discharge capacity degradation during early cycles often leads to inaccuracies due to the influence of various factors (current, voltage, temperature, time) and sudden changes in degradation trends [35, 48].
Data-driven models predict the RUL of LIBs by analyzing data during current cycles. Techniques like dual-input Deep Neural Networks (DNN) [51], 1D Convolutional Neural Networks (1DCNN) [27], Dense layers [24], Long Short-Term Memory (LSTM) networks [17], and Echo State Networks (ESN) [23] have shown superior performance.
@@
Accurately predicting the RUL of LIBs remains challenging due to the varying and volatile degradation of features across cycles, which necessitates a highly adaptable prediction model. The specific challenges will be discussed in the next section.
@@ 1.2 Problem Statement
As summarized in Table 1, current state-of-the-art studies categorize RUL prediction methods into two primary approaches: historical data-independent methods, which estimate the current RUL based on present cycling data and a few preceding cycles, and historical data-dependent methods, which leverage early-cycle data to predict the battery’s full lifespan.
@@
Many existing methods predict RUL based on estimated maximum discharge capacity thresholds, representing the remaining life as a percentage. However, expressing the output as the number of remaining charge-discharge cycles provides clearer.
Small datasets, like the Oxford Battery dataset (13 cells) and NASA battery datasets (4–34 cells) limit model robustness in real-world failure prediction. In contrast, large datasets such as the TRI dataset (124 cells, fast-charging) and the LHP dataset (77 cells, diverse discharge) provide extensive charge-discharge scenarios, making them well-suited for both training and testing of data-independent models.
Signal preprocessing can be limited by high dimensionality, especially with variational decomposition methods like EMD and VMD, which preserve or expand the original signal size. In contrast, statistical feature extraction methods—such as mean and standard deviation—offer a low-dimensional, efficient alternative that captures essential characteristics, making them ideal for real-time and high-accuracy industrial applications.
Model-based and hybrid approaches typically rely on early-cycle data for RUL prediction, yet each battery exhibits unique degradation patterns over its lifespan requiring adaptive data-driven strategies. Moreover, Domain adaptation (DA) techniques such as domain-adversarial neural networks (DANN) [14, 52] and Generative Adversarial Networks (GANs) [16] offer effective solutions for adapting to other source’s degradation patterns to improve RUL prediction in the target domain. To address these challenges, our proposed approach will be represented in the next section.
@@ 1.3 Main Contribution
The main contributions of this work are summarized as follows:
  • We propose a historical data-independent RUL prediction framework for lithium-ion batteries that relies solely on present and recent cycling data, eliminating the need for early-cycle information. This makes the approach suitable for real-time applications and scenarios with incomplete historical records.
  • Our model integrates advanced deep learning components—including LSTM, Multihead Attention, and NODE blocks—as a powerful feature extractor, along with linear layers in the predictors, enabling the capture of both temporal and dynamic battery behaviors.
  • We introduce a domain adaptation strategy that combines two predictors using trainable trade-off parameters and a Maximum Mean Discrepancy (MMD)-based loss to learn domain-invariant features, enhancing transferability from source to target domains.
  • The framework includes a lightweight yet robust preprocessing pipeline-noise reduction, statistical feature extraction (e.g., mean, standard deviation), and normalization—to improve signal quality and reduce dimensionality for efficient, real-time prediction.
  • Extensive evaluations on the two largest publicly available datasets of A123 APR18650M1A cells [44, 51], covering diverse charging and discharging conditions, validate the superior performance and practical applicability of our approach in real-world battery health management.
@@ 2 Preliminaries
This section presents an overview of the key components of the prediction model architecture, including LSTM, Multihead Attention, and Neural Ordinary Differential Equation (NODE) [6] blocks.
2.1 LSTM [17]:
A recurrent architecture designed to overcome the vanishing gradient problem by introducing gating mechanisms. Its operations are defined by:
@@ 2.2 Multihead Attention [47]:
2.3 NODE [6]:
A framework that extends deep learning architecture by modeling continuous-time dynamics instead of discrete transformations between layers. In NODE, the evolution of a hidden state $h(t)$ is governed by an ordinary differential equation (ODE):
@@ 2.3 NODE [6]:
where $f$ is a neural network parameterized by $\theta$. The final state $h(t)$ is obtained by solving this ODE over a time interval, which provides a flexible and memory-efficient representation.
@@ 3.1 Overall architecture
Figure 1: The overall RUL prediction process for Lithium-ion battery cells.
Figure 1 illustrates the RUL prediction process for Lithium-ion battery cells. In the data collection phase, Lithium Iron Phosphate (LFP)/graphite cells are monitored to capture voltage, current, and capacity signals for each individual charge-discharge cycle.
Regarding cycle life degradation, the cycle life of a battery is defined as the total number of charge-discharge cycles from the Beginning of Life (BOL) to the End of Life (EOL). The EOL is typically identified when the battery’s maximum capacity degrades to 70% [41] or 80% [35] of its nominal capacity. The RUL, expressed in terms of the remaining number of cycles, is computed as:
@@
where Aging cycles refers to the number of cycles the battery has already undergone.
In the signal preprocessing phase, the raw signals—voltage, current, and capacity from the most recent charge-discharge cycles—are passed through a noise-reduction filter named median filter [25] to smooth out sudden peaks. The filtered signals are then processed using feature extraction methods, including mean, standard deviation (Std), minimum (Min), maximum (Max), variance (Var), and median (Med) [12, 45]. The extracted features for each cycle are represented as
$X^{i}=\left[x_{\text{current}}^{i},x_{\text{voltage}}^{i},x_{\text{capacity}}^{i}\right],\quad X^{i}\in\mathbb{R}^{3\times 6},$
where the 3 rows correspond to the three signal types (voltage, current, and capacity), and the 6 columns represent the extracted features (Mean, Std, Min, Max, Var, Med). Each input sample to the prediction model consists of 10 selected cycles, uniformly sampled from a 30-cycle window (i.e., one cycle every three cycles) [36]. The input sample is represented as
$\mathbf{X^{i}}=\left[X^{1},X^{2},\dots,X^{10}\right],\quad\mathbf{X}\in\mathbb{R}^{10\times 3\times 6}.$
Thus, the shape of the total target input data after the feature extraction step becomes
$\mathbf{X^{T}}\in\mathbb{R}^{N\times 10\times 3\times 6},$
where $N$ denotes the number of samples. During normalization, a MinMaxScaler is fitted and applied to scale each feature across all time steps and samples between 0 and 1.
In the prediction phase, the RUL prediction model, named HybridoNet-Adapt, maps the target input $\mathbf{X^{T}}$ to the predicted RUL $\mathbf{Y^{T}}\in\mathbb{R}^{N\times 1}$. The details of the proposed RUL prediction model are presented in the following section.
3.2 HybridoNet-Adapt: A Proposed RUL Prediction model with Novel Domain Adaptation
Figure 2: Architecture of the proposed HybridoNet-Adapt framework for RUL prediction with domain adaptation.
As shown in Figure 2, HybridoNet-Adapt is composed of four key components: the source predictor $G^{S}_{Y}$, the target predictor $G^{T}_{Y}$, and the feature extractor $G_{F}$, which is equipped with a DA technique to bridge the distribution gap between the source and target domains.
The feature extractor integrates a LSTM (Section 2.1), a Multihead Attention mechanism (Section 2.2), and a NODE block (Section 2.3). The NODE block models the hidden state $h(t)$, which evolves continuously over time according to the following ODE: $\frac{dh(t)}{dt}=f(h(t),t,\theta),$ where $h(t)$ denotes the hidden state at time $t$, $f$ is a trainable function parameterized by $\theta$, and $t$ represents the continuous time variable. In our implementation, $f$ is realized as a single linear layer to strike a balance between performance and computational efficiency. The initial condition for the NODE block is given by $h(t_{0})$, and the final transformed state $h(t_{1})$ is obtained by solving the ODE over the time interval $[t_{0},t_{1}]$:
@@ 3.2 HybridoNet-Adapt: A Proposed RUL Prediction model with Novel Domain Adaptation
In our experiments, the time bounds are set to $t_{0}=0$ and $t_{1}=1$, based on empirical results (see Figure 8). The function $h(t)$ thus represents the dynamic trajectory of the hidden state under continuous transformation, enabling the model to capture more nuanced temporal dependencies than discrete-time layers alone.
The source and target predictor modules share the same architecture, each consists of three linear layers, each linear layer followed by ReLU [7] activation, 1D Batch Normalization [22], and Dropout [19], respectively, with a sigmoid activation [53] layer at the end.
The target and source predictions in HybridoNet-Adapt are respectively computed as follows:
@@
$\hat{Y}^{S}_{i}(X_{i})$ $=G^{S}_{Y}\big{(}G_{F}(X_{i})\big{)},$ (2)
where $\theta^{S}$ and $\theta^{T}$ are learnable trade-off parameters that balance the contributions from the source and target predictors. The outputs $\hat{Y}^{T}_{i}$ and $\hat{Y}^{S}_{i}$ denote the target and source predictions, respectively.
To optimize the model, we employ a domain adaptation strategy that combines two loss functions: the mean squared error (MSE) [8] loss, $\mathcal{L}_{\text{MSE}}$, used for regression targets $\hat{Y}^{T}_{i}$ and $\hat{Y}^{S}_{i}$; and the maximum mean discrepancy (MMD) [4] loss, $\mathcal{L}_{\text{MMD}}$, which encourages alignment between the feature distributions extracted from the source and target domains. The total loss function is defined as:
@@
$+\lambda\,\mathcal{L}_{\text{MMD}}\big{(}G_{F}(X^{S}_{i}),G_{F}(X^{T}_{i})\big{)},$
where $X^{S}_{i}$ and $X^{T}_{i}$ are the input samples from the source and target domains, respectively, and $Y^{S}_{i}$ and $Y^{T}_{i}$ are their corresponding RUL labels. The hyperparameter $\lambda$ controls the weight of the MMD loss in the overall objective.
The MMD loss quantifies the discrepancy between the distributions of source and target feature embeddings. Given extracted feature sets $F^{S}_{i}$ from the source domain $G_{F}(X^{S}_{i})$ and $F^{T}_{j}$ from the target domain $G_{F}(X^{T}_{j})$, it is defined as:
$\mathcal{L}_{\text{MMD}}(F^{S}_{i},F^{T}_{j})$ $=\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}k(F^{S}_{i},F^{S}_{j})+\frac{1}{m^{2}}\sum_{i=1}^{m}\sum_{j=1}^{m}k(F^{T}_{i},F^{T}_{j})$
$\quad-\frac{2}{nm}\sum_{i=1}^{n}\sum_{j=1}^{m}k(F^{S}_{i},F^{T}_{j}),$ (3)
where $k(\cdot,\cdot)$ is a kernel function, commonly chosen as the Gaussian kernel: $k(x,y)=\exp\left(-\frac{\|x-y\|^{2}}{2\sigma^{2}}\right),$ with $\sigma$ as the kernel bandwidth parameter. $n$ and $m$ denote the number of training samples from source and target domains, respectively.
The MSE loss is used to optimize the regression outputs by penalizing the squared differences between predicted and ground truth values. It is defined as:
$\mathcal{L}_{\text{MSE}}(\hat{Y_{i}},Y_{i})=\frac{1}{u}\sum_{i=1}^{u}\left(\hat{Y_{i}}-Y_{i}\right)^{2},$
where $\hat{Y}_{i}$ denotes the predicted value, and $Y_{i}$ is the corresponding label for the RUL. $u$ is the number of training samples.
The hidden dimension of both LSTM and Multihead Attention is set to 64, with 2 LSTM layers. The hidden dimensions of the linear layers in the source and target predictors are $[128,64,32,1]$, with a dropout rate of 0.1. These values are determined based on the experimental results presented in Section 4.6.
In the following section, a series of experiments are conducted to identify the optimal configuration of HybridoNet-Adapt and to demonstrate its superiority over state-of-the-art methods. For comparison purposes, a supervised baseline model named HybridoNet is constructed, consisting of the target predictor $G^{T}_{Y}$ and the feature extractor $G_{F}$. This model is trained using the MSE loss function. By comparing its performance with that of HybridoNet-Adapt, we highlight the performance improvements achieved through the incorporation of our proposed Domain Adaptation technique.